# Eliminating Warping Shakes for Unsupervised Online Video Stitching

### v1

@@ 0.B.2 Loss Function
The alignment loss and distortion loss for the spatial/temporal warp model are also similar to UDIS++ [46]. One can refer to [46, 45] for more details. For the convenience of readers, we paraphrase their definitions here again.
@@
where is the reference/target frame, is an all-one matrix with the same size as , and is a constant to balance different transformations.
@@ Appendix 0.C Warp Smoothing Model
Appendix 0.D Evaluation Metric
In the beginning, we evaluate the distortion and stability performance with the metrics that are widely used in video stabilization [36, 34, 64]. These traditional metrics try to estimate the spatial transformation (homography or affine) between adjacent frames from keypoint correspondences. However, the estimated point correspondences are unreliable in our challenging testing cases (e.g., low texture or low light). In addition, as described in Nie et al.’s video stitching [48], the metric in the frequency domain (i.e., the stability score in [36]) are not reliable sometimes as the trajectory signals are usually very short and of different lengths. Therefore, we adopt the more intuitive indicators (i.e., the distortion loss and smoothness loss) to describe the distortion and stability performance. Considering the trajectories and mesh resolutions are defined differently in different solutions, we only apply the metrics in our ablation studies to show the effectiveness of each module.
@@ Appendix 0.D Evaluation Metric
| 6 | StabStitch | 23.39 | 19.69 | 18.53 | 134.61 | 49.05 |
Appendix 0.E More Experiment

### v2

@@ 0.B.2 Loss Function
The alignment loss and distortion loss for the spatial/temporal warp model are also similar to UDIS++ [47]. One can refer to [47][46] for more details. For the convenience of readers, we paraphrase their definitions here again.
@@
where / is the reference/target frame, is an all-one matrix with the same size as , and is a constant to balance different transformations.
@@ Appendix 0.C Warp Smoothing Model
Appendix 0.D Dataset
The videos in our dataset consist of three parts: some original videos from [64], some stable videos from [57], and our captured videos. These videos are captured with arbitrary and irregular motion trajectories. Therefore, we leverage them to simulate two videos from different perspectives. Specifically, we collect the video pair from different timestamps (e.g., one video is from the original video, and the other video is captured after a random delay time). After that, we crop the video frames to simulate an appropriate overlapping rate in stitching. Considering the videos are collected from different timestamps, we further filter out the videos with obvious moving objects. Finally, we get over 100 video pairs and demonstrate the distribution of video duration in Fig. 9.
Appendix 0.E Evaluation Metric
Please note that in the comparative experiments, we only adopt the alignment score because different methods define different warp representations and camera trajectories. Thus we only apply the last two metrics to our ablation studies to show the effectiveness of each module.
In the beginning, we evaluate the distortion and stability performance with the metrics that are widely used in video stabilization [36][34][66]. These traditional metrics try to estimate the spatial transformation (homography or affine) between adjacent frames from keypoint correspondences. However, the estimated point correspondences are unreliable in our challenging testing cases (e.g., low texture or low light). In addition, as described in Nie et al.’s video stitching [49], the metric in the frequency domain (i.e., the stability score in [36]) are not reliable sometimes as the trajectory signals are usually very short and of different lengths. Therefore, we adopt the more intuitive indicators (i.e., the distortion loss and smoothness loss) to describe the distortion and stability performance.
@@ Appendix 0.E Evaluation Metric
| 6 | StabStitch | 23.18 | 19.53 | 18.67 | 133.59 | 48.74 |
Appendix 0.F More Experiment